This week a factory is producing 50 units of a particular commodity, and the amount being

produced is increasing at the rate of 2 units per week. If 𝐢(π‘₯) dollars is the total cost of producing

π‘₯ units and 𝐢(π‘₯) = 0.08π‘₯^3 βˆ’ π‘₯^2 + 10π‘₯ + 4 , find the current rate at which the production cost is increasing.








1
Expert's answer
2020-12-07T18:32:01-0500

"C(x) = 0.08x^3 - x^2 + 10x + 48\\

\frac{\mathrm{d}C}{\mathrm{d}x} = 0.24x^2 - 2x + 10\\

\begin{aligned}
\frac{\mathrm{d}C}{\mathrm{d}x}\biggr\vert_{x = 2} &= 0.24(2)^2 - 2(2) + 10
\\&= 0.24\times 4 - 4 + 10
\\&= 0.96 - 4 + 10 = 6.96
\end{aligned}\\

 
\textsf{Therefore, the current rate at}\\
\textsf{which the production cost}\\
\textsf{is increasing is}\, \, 6.96 \\
\textsf{dollars per units.}"